### v2

# Line of Sight Bias in Dark Matter Inferences from Galaxy Cluster Mergers

Abstract
In collisions of galaxy clusters, the lack of displacement between dark matter and galaxies suggests that the dark matter scattering depth is small. This yields an upper limit on the dark matter cross section if the dark matter column density is known. We investigate a bias in such constraints: the measured column density (along the line of sight, using gravitational lensing) is lower than that experienced by a dark matter particle, as follows. Dark matter halos are triaxial and generally collide along their major axes, yielding a high scattering column density—but the merger is obvious only to observers whose line of sight is nearly perpendicular to that axis, yielding a low observed column density. We trace lines of sight through merging halos from the BigMDPL n-body simulation, both with and without mock observational effects. We find that a hypothetical skewer through the center of the halo along the merger axis (more precisely, along the current separation vector of the two halos) has nearly twice the column density of a typical line of sight. With weak lensing measurements, which involve some spatial averaging, this ratio is reduced to 1.25, suggesting that existing constraints on the scattering cross section are biased high by about 25%.
Galaxy clusters (584); dark matter (353)
1 Introduction
Dark matter (DM) is the dominant form of matter in the universe, but little is known about its particle properties. One key question is to what extent, if any, DM particles can scatter off each other. The astrophysical effects of such scattering, also known as self-interacting dark matter (SIDM), were first discussed by Spergel & Steinhardt (2000), who also presented rough upper limits on the scattering cross section per unit mass $\frac{\sigma_{\rm DM}}{m}$ based on those effects. In many particle models, $\frac{\sigma_{\rm DM}}{m}$ will be velocity-dependent, so constraints from different astrophysical environments should be considered in the context of the typical particle velocity in that environment (Kaplinghat et al., 2016). The highest velocities probed are in the mergers of galaxy clusters, with relative speeds of $\sim 3000$ km/s (Markevitch et al., 2004; Randall et al., 2008; Harvey et al., 2015; Robertson et al., 2017; Wittman et al., 2018b).
Markevitch et al. (2004) identified the Bullet cluster as a post-pericenter snapshot of a nearly head-on collision between two galaxy clusters. A hallmark of recent pericenter passage is a substantial separation between gas and galaxies (later called dissociation by Dawson et al. (2011)) due to momentum exchange between the gas distributions around the time of pericenter. (Markevitch et al., 2004) argued that the lack of separation between DM and galaxies implied that the DM scattering depth $\tau=\frac{\sigma_{\rm DM}}{m}\Sigma<1$, where $\Sigma$ is the surface mass density encountered by a DM particle of one halo as it passes through the other halo. Hence $\Sigma^{-1}$ yields an upper limit on $\frac{\sigma_{\rm DM}}{m}$ if no significant galaxy-DM separation is observed. Harvey et al. (2014) generalized this argument to yield an estimate of $\frac{\sigma_{\rm DM}}{m}$ if a separation is observed.
This paper explores a bias in this method. In practice $\Sigma$ is measured along the line of sight (LOS), but for purposes of the DM constraint it should be measured along the merger trajectory, or along the current separation vector (CSV) as a proxy. Because halos are triaxial (Jing & Suto, 2002) and preferentially aligned with neighboring halos (Binggeli, 1982; Plionis, 1994; Kasun & Evrard, 2005; Smargon et al., 2012), the surface mass density probed along the CSV, $\Sigma_{\rm CSV}$, will tend to be larger than the surface mass density probed along a random LOS, $\Sigma_{\rm los}$. Furthermore, the merging systems used for these constraints tend to be those in which the LOS is perpendicular to the CSV, such that physical separations would be visible as angular separations. This implies that $\Sigma_{\rm los}$ is measured along a triaxial halo’s intermediate or minor axis, biasing it low compared to a random LOS. If the ratio $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}<1$, then the DM scattering depth is underestimated and $\frac{\sigma_{\rm DM}}{m}$ is overestimated. This paper examines dark matter-only n-body simulations to quantify the factor by which $\frac{\sigma_{\rm DM}}{m}$ is overestimated.
In §2 we describe our methods, in §3 we present the results, and in §4 we discuss the impact of the results.
2 Methods
We approximate each halo as a triaxial ellipsoid, meaning that its density $\rho$ is a function only of the scaled radial coordinate $r\equiv\sqrt{(\frac{x}{a})^{2}+(\frac{y}{b})^{2}+(\frac{z}{c})^{2}}$ where $a$, $b$, and $c$ denote the semiaxis lengths ($a>b>c$). We assume for notational simplicity here a Cartesian coordinate system aligned with the halo principal axes; in practice, this can be achieved by a rotation from the coordinate system of the simulation in which the halo is embedded.
2.1 Central Skewers
The surface mass density $\Sigma$ along any skewer piercing the center of the ellipsoid can be expressed as the skewer distance across the ellipsoid times an integral $\int_{0}^{\infty}\rho(r)dr$ that depends only on the density profile. Therefore the ratio $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ cancels the latter factor and involves only the ratio of distances across the ellipsoid. In this subsection we argue that this simple ratio is a useful thinking tool because it provides a bound on the size of the effect when integrated over larger surface areas, independent of the density profile or observational technique. Later in the paper, we perform more realistic estimates with mock weak gravitational lensing fits, which inherently average over a patch of sky.
Figure 1: Mean surface mass density along the minor axis, $\bar{\Sigma}_{\rm minor}$, as a fraction of that along the major axis, $\bar{\Sigma}_{\rm major}$, as a function of projected radius of the area enclosed, for a halo with minor axis ratio $\frac{c}{a}=0.4$. The central skewer prediction of 0.4 provides a lower bound independent of density profile.
Consider a LOS parallel to the major ($x$) axis, but displaced in $z$: Taylor expanding the density about $z=0$ produces a term proportional to $\frac{\rho^{\prime}(r)}{c^{2}}$. In contrast, expanding a minor-axis LOS about $x=0$ produces a term proportional to $\frac{\rho^{\prime}(r)}{a^{2}}$. Therefore, the density drops more quickly for LOS displaced from the major axis; including these non-central effects over a finite patch of sky will decrease $\bar{\Sigma}_{\rm major}$ faster than it decreases $\bar{\Sigma}_{\rm minor}$, where the bars indicate area averaging. Hence, skewers through the center provide a useful lower bound on the ratio of mean surface mass densities averaged over finite areas. We confirm this by using the formalism of Heyrovský & Karamazov (2024) to find $\frac{\bar{\Sigma}_{\rm minor}}{\bar{\Sigma}_{\rm major}}$ averaged over circles centered on the principal axes, for a variety of density profiles and circle sizes. Figure 1 shows the result for a halo with $\frac{b}{a}=0.6$ and $\frac{c}{a}=0.4$, for three profiles: NFW (Navarro et al., 1997), Einasto (Einasto, 1965), and Gaussian. The scale radius of the Einasto profile was chosen to match the NFW. The Gaussian profile is not physically motivated, but illustrates the behavior of an entirely different profile. With $\frac{c}{a}=0.4$, the central skewer argument predicts $\frac{\Sigma_{\rm minor}}{\Sigma_{\rm major}}=0.4$. Near the center, the Gaussian density profile is close to flat so the ratio remains close to 0.4 even when averaged over a substantial area. The other profiles are not flat near the center so $\frac{\bar{\Sigma}_{\rm minor}}{\bar{\Sigma}_{\rm major}}$ increases more rapidly with averaging radius. In all cases the central skewer ratio of 0.4 provides a lower bound, and overestimates the departure from unity by less than a factor of two.
Having established the utility of the central skewers, we now proceed to geometric definitions that will be used for both the skewer estimates and the mock weak lensing. At the end of this section we transition to describing the weak lensing simulations.
2.2 Geometric definitions
Figure 2 illustrates the geometry used in this paper. The black arrow defines the first principal axis (major axis) of the triaxial halo, while the red arrow shows an illustrative CSV (the other halo involved in the merger is not shown). Halos tend to be aligned with their neighbors, so the angle $\psi$ is generally small as shown. The LOS is drawn at a larger angle $\alpha$ from the separation vector, because lines of sight near the separation vector would not identify a merger as dissociative. There are techniques for constraining $\alpha$ (e.g., Wittman et al., 2018a) but these do not constrain the relationship between the LOS and the halo principal axes. Hence, the LOS could be anywhere along a cone with opening angle $\alpha$. Figure 2 shows two possible LOS along this cone, illustrating that one involves a considerably longer path through the halo hence a larger $\Sigma_{\rm los}$.
Figure 2: Geometric view. $\psi$ separates the halo major axis and the current separation vector to the other halo in the merger (not shown here). Lines of sight (LOS) fall along a cone defined by the separation vector and $\alpha$. At fixed $\psi$ and $\alpha$, LOS can vary substantially in column density.
We calculate the distance through the halo along the CSV, and along all possible LOS (at a given $\alpha$) to obtain a distribution for $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ given a specific halo, merger, and $\alpha$. We can then marginalize over many halos and mergers to obtain a distribution for $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ given $\alpha$. The selection of mergers could be tailored to match a specific observed system as in Wittman (2019), but this paper derives results for mergers in general.
2.3 Simulation inputs and calculations
[…]
We use the catalog of halo pairs assembled in Wittman et al. (2018a) and Wittman (2019). Briefly, this is derived from the BigMDPL Rockstar (Behroozi et al., 2013) halo catalog by finding halo pairs that have passed exactly one pericenter (within 300 kpc). This is in principle a different catalog at each snapshot, but we focus on one illustrative snapshot (number 74, redshift $0.1058$). We discarded halo pairs with current separations <0.25<0.25 Mpc, as these were too overlapping for Rockstar to derive robust shape parameters. Observed mergers with current separations <0.25<0.25 Mpc are also rare because they are observationally difficult to disentangle.
We use the axis1_[xyz] attributes (square brackets denote use of each enclosed letter) to determine the orientation of the halo major axis. The orientations of the second and third principal axis are not documented, so we marginalize over them as described below. We used the axisratio_[23]_1 attributes to quantify the axis ratios. For a given halo, we first compute the angle $\psi$ between the halo’s major axis and the CSV given by the halo-pair catalog; see Figure 2. We then calculate the skewer lengths (hence $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$) for a variety of LOS, and finally group the skewer results by the angle $\alpha$ between the CSV and the LOS.
The principal axes in the BigMDPL database are determined using the method of Allgood et al. (2006), which starts with a spherical region of radius of $R_{\rm vir}$ and calculates a reduced inertia tensor considering the particles within this region. Then, they update the region to an ellipsoidal region using the obtained axis ratios, keeping the semimajor axis fixed at $R_{\rm vir}$, and redo the calculation, iterating this process until it converges. This is a widely adopted robust way to determine axis ratios, and Bett (2012) showed that the median axis ratios produced this way are similar (within 0.05) to those yielded by a simple inertia tensor. We found that the median $\frac{c}{a}$ for BigMDPL halos above $0.6\times 10^{14}h^{-1}M_{\odot}$ is comparable to but slightly lower than those found in other studies: 0.46 vs. 0.45–0.50 for a similar mass range in Tenneti et al. (2015) and 0.55 for the ‘reduced iterated’ method in Bett (2012). The greater elongation of BigMDPL halos implies that we would find slightly smaller LOS bias effects in these other simulations. Still, it is not clear that BigMDPL halos are unduly elongated for our purposes given that halos are more elongated at higher density thresholds (Jing & Suto, 2002) where DM scattering effects will be most important. Our median $\frac{c}{a}$ of 0.46 agrees well with the median $\frac{c}{a}$ found in Jing & Suto (2002) for the high density threshold corresponding to the physical radius of 150 kpc we adopt below for averaging the surface mass density, even though the BigMDPL shapes were not explicitly evaluated at these high densities.
We checked the reliability of these axis ratios in a few ways. First, the BigMDPL database includes a second set of axis ratios, axisratio_[23]_1_500c, which are measured at a smaller radius than $0.3R_{\rm vir}$. We found that the statistics were quite similar, with median axis ratios differing only by $\approx 0.01$. Second, we found that in BigMDPL binary mergers, the axis ratios for the lower-mass partners match those of isolated halos (median $\frac{c}{a}=0.46$), but the more massive (’main’) partner is more elongated ($\frac{c}{a}=0.35$). This difference cannot be explained by particle misattribution near pericenter, which would presumably affect the lower-mass halo more. At least part of this difference may be the known effect that higher-mass halos are more elongated.
2.4 Weak Lensing
Inferences made via weak lensing average over a patch of sky and thus, unlike the central skewers, depend on the density profile. However, Figure 1 shows little difference between profiles commonly used to fit clusters. We therefore adopt the more commonly used NFW profile. We use the cataloged BigMDPL halo parameters m200c, rs, axisratio_2_1, and axisratio_3_1 to create a triaxial NFW model and fit it as follows. The workflow is summarized in Figure 3. Following Corless & King (2008), we use the spherical scale radius rs and the axis ratios to define a triaxial scale radius $R_{s}$, then use their Equation 7 to recover the triaxial $R_{200}$ given m200c. The ratio $\frac{R_{200}}{R_{s}}$ then forms a triaxial concentration which yields a scale density $\rho_{s}$. We then use the formalism of Heyrovský & Karamazov (2024) to obtain the true $\bar{\Sigma}_{\rm CSV}$ of the triaxial halo, which we store for reference. For this calculation, we use a fixed physical radius of 150 kpc as was used in the seminal dark matter constraint of Markevitch et al. (2004), although we vary this radius later to probe the robustness of the results.
Figure 3: Summary of the workflow for establishing a triaxial model halo and extracting inferences from mock observations.
We then choose a LOS and use the expressions in Heyrovský & Karamazov (2024) to generate a reduced shear field for that LOS. We discard any “observations” of shear in a 30′′ radius area around the center of the cluster, where cluster galaxies would prevent the detection of many background galaxies, and where the weak lensing approximation breaks down for massive clusters. To set the size of the simulated field, we tested a series of sizes with spherical halos and found that at 20 and larger diameter, the fitted mass converges to the true mass. We opted to use a 30 diameter field for the main run over all halos and all LOS.
To simulate the process of estimating $\bar{\Sigma}$ as it would typically be done in an observational paper, we then fit the mock shear field with a spherical NFW model following the Child et al. (2018) mass-concentration ($M$-$C$) relation, as weak lensing data are typically not informative enough to infer concentration or shape parameters. We fit with two free parameters, $M$ and $C$, which are linked by the $M$-$C$ relation but allowed to vary according to the scatter found by Child et al. (2018). In order to focus on the specific bias we are highlighting, we did not explore the effect of adding noise to the shear field. In that context, the central sky coordinates of the NFW halo were always well fit when allowed to float, so for efficiency in the main run we fixed those coordinates, leaving $M$ and $C$ as the only two free parameters.
For each halo and LOS, we use the best-fit spherical model to infer $\bar{\Sigma}_{\rm los}$ over a 150 kpc radius patch and normalize it by the known $\bar{\Sigma}_{\rm CSV}$ for that halo. In §3 we explore how $\frac{\bar{\Sigma}_{\rm los}}{\bar{\Sigma}_{\rm CSV}}$ varies as a function of viewing angle and halo properties.
3 Results
3.1 Halo-merger alignment
Figure 4 shows the aligment between the CSV and each halo’s major axis, in terms of a histogram of the angle $\psi$. The two axes show a remarkable tendency for alignment (small $\psi$), in contrast to the dashed curve which shows the distribution expected for CSVs randomly distributed on a sphere. Assuming the clusters fell toward each other along a connecting filament, this is consistent with work finding that cluster-scale halos are aligned along larger-scale filaments, which aligns them with neighboring halos (Kasun & Evrard, 2005).
Figure 4: Histogram of the angle $\psi$ between the merger axis and each halo’s major axis. The two axes show a remarkable tendency for alignment (small $\psi$), in contrast to the dashed curve which shows the distribution expected for $\psi$ distributed randomly on a sphere.
For binary merging clusters specifically, we would expect good halo-halo alignment because both halos are falling toward each other along a filament with which they are each aligned. We are not aware of empirical measurements of this effect, which may be difficult given overlapping halos. However, halo-halo alignment in binary mergers can be inferred from two empirical results: the orientation of a brightest cluster galaxy (BCG) is a proxy for the orientation of its host cluster (West et al., 2017), and BCGs in merging clusters are known to align with the CSV (Wittman et al., 2019). Although we do not attempt a direct comparison because those works involve 2-D measurements of galaxies rather than 3-D measurements of clusters, the alignment seen in Figure 4 is about as strong as seen in those works.
3.2 Skewer surface mass density
Illustrative halos. Figure 5, top panel, illustrates how $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ varies with $\alpha$ for a very well aligned and very elongated halo ($\psi=1.5^{\circ}$, $\frac{b}{a}=0.337$, $\frac{c}{a}=0.259$). Results are shown as violin plots for values of $\alpha$ in 5 increments. At $\alpha=0$, $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}=1$ by definition and only one LOS is sampled so no violin appears. As $\alpha$ increases, the LOS pierces the halo closer to the intermediate or minor axis, so $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ falls below unity. Furthermore, at each $\alpha>0$ there are multiple LOS, yielding a range of $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ depending on their proximity to the halo minor or intermediate axis. Observed systems are typically identified as mergers and used for DM constraints only if $\alpha>45^{\circ}$, so this halo suggests that $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}\approx 0.3$. In other words, the surface mass density encountered by DM particles through the pericenter passage in this rather extreme case is about three times that measured along the LOS, so the resultant DM constraint would be biased high by a factor of three.
Figure 5: Skewer bias factor $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ as a function of viewing angle $\alpha$ for two halos with similar axis ratios. Top: halo 13410052637 is well aligned with its merger axis ($\psi=1.5^{\circ}$), such that lines of sight perpendicular to the merger axis (large $\alpha$) encounter systematically low column density. Bottom: halo 13582533817 is less well aligned ($\psi=18^{\circ}$), leading to more mixed results but still with bias at large $\alpha$. In each case the violin shows the distribution, which the open circle marks the median and the black band marks the middle quartiles.
Figure 5, bottom panel, shows an equally elongated halo with a more typical $\psi=18^{\circ}$ misalignment from the CSV. Therefore the maximum $\Sigma_{\rm los}$ is encountered by one LOS at $\alpha=18^{\circ}$: the LOS that happens to align with the major axis. Other LOS at $\alpha=18^{\circ}$ encounter a range of lower column densities. As one proceeds to higher $\alpha$ the LOS are sampling the intermediate and minor axes—but to a lesser degree than with a well-aligned halo, hence $\Sigma_{\rm los}$ does not drop as far. As a reminder, these are highly elongated halos chosen to illustrate the dependence of $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ on halo alignment $\psi$ and viewing angle $\alpha$. We now proceed to marginalize over all merging halos.
All halos. A few halo pairs provided remarkably large or small bias factors, for example when one halo’s major axis points at the other’s minor axis in a “T” shape. We suspect that these extreme cases are artifacts of the halo finder having difficulty separating overlapping halos. To reduce clutter from these possibly unphysical cases, in Figure 6 we show the median and interquartile range (IQR) of the bias factor distribution after marginalizing over all halos in the relevant halo pairs. Again, considering that observed cases will have $\alpha>45^{\circ}$, the surface mass density encountered by DM particles through the pericenter passage is nearly twice that measured along the LOS. Given that the central skewer estimates are a lower limit on $\frac{\bar{\Sigma}_{\rm los}}{\bar{\Sigma}_{\rm CSV}}$, we expect that resulting DM constraints would be biased high, but by less than a factor of two.
Figure 6: Median (circles) and interquartile range (bars) of the skewer bias factor marginalized over all halos, as a function of viewing angle $\alpha$. For typical observed cases $\alpha>45^{\circ}$ so the central surface mass density along the LOS is about half that encountered by DM particles along the separation vector. The gray curves illustrate the median behavior of each individual halo (medianized over various LOS at fixed $\alpha$).
Effect of mass. Figure 7 shows how these trends differ across halo mass quartiles; for clarity only the median effect at each $\alpha$ and mass bin is shown. Lower-mass halos are rounder (in BigMDPL as well as in other works, e.g. Allgood et al., 2006; Henson et al., 2016) so the effects of LOS are less marked at lower mass.
Figure 7: Median skewer bias factor $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ as a function of viewing angle $\alpha$, marginalized over all halos in a given mass quartile. Lower-mass halos show reduced LOS effects because they are rounder.
3.3 Mock weak lensing results
We show the WL results as a function of $\alpha$ in Figure 8. Each halo is shown as a gray curve, with the medians and interquartile ranges overlaid. As expected, the results are not as extreme as the central skewer calculation: at large $\alpha$ the bias is about 0.8. This LOS bias happens to be of similar magnitude to that found by Euclid Collaboration et al. (2024) for total mass. They performed mock WL fits along the principal axes of clusters in the Three Hundred simulation (Cui et al., 2018) and found that orientations along the (major, intermediate, minor) axes bias the inference of the true mass by factors of about (1.2, 0.85, 0.75). [^2] The Three Hundred halos include baryons so should be slightly rounder than our BigMDPL halos which are dark matter only. However, this is a small effect; see §4. However, this is a numerical coincidence because our calculation is for surface mass density within 150 kpc projected radius rather than mass, and we show below that the result is sensitive to that particular choice.
Figure 8: Weak lensing bias factor as a function of viewing angle $\alpha$, marginalized over all halos. The circles indicate medians and the bars indicate the interquartile range. The gray curves illustrate the median behavior of each individual halo (medianized over various LOS at fixed $\alpha$). The trend is similar to that seen for the skewer estimate, but the bias factor is less extreme.
Note one distinction between the construction of the skewer and WL plots: for the skewers, the ratio $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ was defined to be unity at $\alpha=0$. For the WL plots, observational effects and modeling choices could cause the inferred $\bar{\Sigma}_{\rm los}$ to differ from the true $\bar{\Sigma}_{\rm CSV}$ even when the LOS is along the CSV. (In fact, Figure 8 shows that the median ratio is slightly above unity at $\alpha=0$.) This choice in plot construction also forces the various halo curves to meet at $\alpha=0$ in the skewer plot (Figure 6) but not the WL plot (Figure 8). This causes an apparent variation in halo-to-halo scatter as a function of $\alpha$ in Figure 6.
Figure 9 shows the same results split by mass quartile. As expected, we find that the higher-mass halos are most sensitive to LOS effects, because they have more extreme axis ratios.
Figure 9: Median weak lensing bias factor as a function of viewing angle $\alpha$, marginalized over all halos in a given mass quartile.
Figure 10: Our fiducial weak lensing bias factor was defined in terms of surface mass density projected within 150 kpc of the halo center. Here we show how the bias factor depends on this choice.
To this point, the weak lensing bias factor has been defined in terms of surface mass density projected within 150 kpc of the halo center. Figure 10 shows the median bias (marginalized over all halos) as a function of $\alpha$ for two other choices of radius, 50 and 300 kpc. There is substantial variation, but mostly at small $\alpha$ which will rarely apply to observations in practice. At large $\alpha$, the variation is about $\pm 0.05$. This is comparable to the variation that might be seen with the use of different halo catalogs as discussed in §2.3.
4 Summary and discussion
We have found that current $\frac{\sigma_{\rm DM}}{m}$ constraints based on the scattering depth argument are on average biased because spherical symmetry is violated: the surface mass density along the LOS is lower than that experienced by DM particles as they pass along the merger axis. Because the surface mass density is inversely related to the inferred cross section, this yields an overestimate of the DM cross section. We cast our results in terms of $\frac{\bar{\Sigma}_{\rm los}}{\bar{\Sigma}_{\rm CSV}}$, which in this argument is the same as the ratio of true to inferred DM cross section. We first introduced a thinking tool—skewers running through triaxial halo centers—that provides a strict lower bound to this ratio, independent of the density profile. In §3.2 we found that this ratio had a median value of 0.5 for typically used LOS, and with large scatter from halo to halo. Our weak lensing simulations (§2.4) show that in practice the ratio is about 0.8 for typically used LOS, and with less scatter. It also shows a modest dependence on the size of the region over which the surface mass density is averaged. Note that this bias is distinct from the overall mass bias studied in Lee et al. (2023), which is smaller and stems from the overlap of two hypothetically spherical halos.
We studied dark matter only (DMO) halos; the presence of baryons will reduce bias by making halos rounder, but only slightly. For example, the simulations of Henson et al. (2016) found that the median minor/major axis ratio shifts from $0.537$ for DMO to $0.576$ with baryons. This is only 40% of the scatter across halos. The intermediate/major axis ratio shifts by a similarly small amount when baryons are added. Still, there remains some uncertainty regarding the size of the effect given that different simulations with the same physics have different median axis ratios: the axis ratios for our BigMDPL halos are nontrivially lower than the DMO value in Henson et al. (2016). A comparable uncertainty is due to the size of the region within which $\bar{\Sigma}$ is measured (Figure 10), so further study is warranted—although some of these uncertainties will be obviated when DM constraints are derived from simulations as described in our final paragraph.
The bias is a function of halo and LOS properties, so more precise estimates could in principle be tailored to each individual observed system. Such tailoring would involve a straightforward extension of the procedure here: rather than counting all halo pairs with equal weight in the marginalization, one should weight by the likelihood of a halo pair matching the observables (subhalo projected separation, mass, and line-of-sight relative velocity).
Most known binary systems were identified due to their clear separation between subclusters on the sky, hence $\alpha$ is not small. This is why we have referred to lines of sight at large $\alpha$ as “typical.” This is likely to remain the case for any cluster amenable to this DM scattering argument; for mergers along the LOS there is no mechanism for measuring a DM offset.
Although our argument was framed in terms of analytical approximations relating $\frac{\sigma_{\rm DM}}{m}$ to scattering depth (Markevitch et al., 2004; Harvey et al., 2014), the same bias applies to staged simulations based on the assumption of spherical symmetry. To overcome this bias, staged simulations should be staged with appropriate halo shapes and alignments, and we urge simulators to do so. We checked the time evolution of the alignment, and we found that main halos are already well aligned at least ${\sim}1$ Gyr before pericenter. Hence one cannot rely on the alignment arising organically from processes (e.g. tidal effects) already present in the staged simulation. A further step would be to go beyond the triaxial approximation and resimulate merging systems found in cosmological simulations. This would incorporate the effect of finer substructure and the fact that axis ratios vary with density (Jing & Suto, 2002). Such simulations would also naturally incorporate the appropriate range of halo alignments and pericenter distances.
We thank the anonymous referee for numerous constructive suggestions. This work was supported by NSF grant number 2308383. The CosmoSim database used in this paper is a service by the Leibniz-Institute for Astrophysics Potsdam (AIP). The MultiDark database was developed in cooperation with the Spanish MultiDark Consolider Project CSD2009-00064. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) and the Partnership for Advanced Supercomputing in Europe (PRACE, www.prace-ri.eu) for funding the MultiDark simulation project by providing computing time on the GCS Supercomputer SuperMUC at Leibniz Supercomputing Centre (LRZ, www.lrz.de).